OpenStudy (anonymous):

Solve 7x - 2y = -3 14x + y = 14

OpenStudy (texaschic101):

do you prefer substitution or elimination ?

OpenStudy (texaschic101):

well....lets just do substitution...it is probably easier. 14x + y = 14 ---> y = -14x + 14 now sub -14x + 14 in for y in the other equation to find x 7x - 2y = -3 7x - 2(-14x + 14) = -3 can you take it from here or do you still need help ?

OpenStudy (anonymous):

is it 14x+4=-6

OpenStudy (anonymous):

im doing the same type of problem texaschic101 i need help whats the next step

OpenStudy (anonymous):

except mine is y=4x+6 and y^2+x^2=40,000

OpenStudy (texaschic101):

I keep getting the wrong answer with substitution, so I am gonna do elimination 7x - 2y = -3 14x + y = 14 --->(2) ------------ 7x - 2y = -3 28x + 2y = 28 (result of multiplying by 2) -------------add 35x = 25 x = 25/35 x = 5/7 now sub 5/7 in for x in either of the equations to find y 7x - 2y = -3 7(5/7) - 2y = -3 35/7 - 2y = -3 5 - 2y = -3 -2y = -3 - 5 -2y = -8 y = 4 check... 14x + y = 14 14(5/7) + 4 = 14 70/7 + 4 = 14 10 + 4 = 14 14 = 14 (correct) x = 5/7 and y = 4

OpenStudy (anonymous):

texaschic101 i need ur help on my problem